Jones Products manufactures and sells to wholesalers approximately 200,000 packages per year of underwater markers at $3.90 per package. Annual costs for the production and sale of this quantity are shown in the table.

Direct materials $ 240,000
Direct labor 60,000
Overhead 192,000
Selling expenses 80,000
Administrative expenses 53,000
Total costs and expenses $ 625,000


A new wholesaler has offered to buy 33,000 packages for $3.40 each. These markers would be marketed under the wholesaler’s name and would not affect Jones Products’s sales through its normal channels. A study of the costs of this additional business reveals the following:

  • Direct materials costs are 100% variable.
  • Per unit direct labor costs for the additional units would be 50% higher than normal because their production would require overtime pay at 1½ times the usual labor rate.
  • 35 percent of the normal annual overhead costs are fixed at any production level from 150,000 to 300,000 units. The remaining 65% of the annual overhead cost is variable with volume.
  • Accepting the new business would involve no additional selling expenses.
  • Accepting the new business would increase administrative expenses by a $2,000 fixed amount.


Required:
Complete the three-column comparative income statement that shows the following (Round your intermediate calculations and per unit cost answers to 2 decimals.)

1. Annual operating income without the special order.
2. Annual operating income received from the new business only.
3. Combined annual operating income from normal business and the new business.
Answer #1

Direct material cost per package = Total direct material costs / 200,000 packages = $240,000 / 200,000 = $1.20

Direct labor cost per package = Total direct labor costs / 200,000 packages = $60,000 / 200,000 = $0.30

Direct labor cost for the special order = Direct labor cost per package + 50% = $0.30 + 50% = $0.45

Fixed overhead = 35% of $192,000 = $67,200.

Variable overhead = 65% of $192,000 = $124,800.

Variable overhead per package = $124,800 / 200,000 packages = $0.62 (rounded off to two decimal places)

S. No. Particulars Annual operating income without special order Annual operating income with new business only Combined Annual operating income
a. Sales Revenue $3.90 * 200,000 = $780,000 $3.40 * 33,000 = $112,200 $892,200
b. Direct materials costs $240,000 $1.20 * 33,000 = $39,600 $279,600
c. Direct labor costs $60,000 $0.45 * 33,000 = $14,850 $74,850
d. Fixed overhead $67,200 0 $67,200
e. Variable overhead $124,800 $0.62 * 33,000 = $20,460 $145,260
f. Selling expenses $80,000 0 $80,000
g. Administrative expenses $53,000 $2,000 $55,000
h. Total Costs and expenses (b+c+d+e+f+g) $625,000 $76,910 $701,910
i. Annual operating income (a-h) $155,000 $35,290 $190,290